# Marginal Metals and Kosterlitz-Thouless Type Phase Transition in Disordered Altermagnets

### v1

@@ Abstract
We introduce disorder into the two AM layers and calculate the tunneling conductances as functions of disorder strength $W$ [Fig. 5(c)]. We find that a substantial conductance difference between $G_{P}$ and $G_{AP}$ persists over a wide range of $W$, extending even to values $W>W_{c}$. This reflects the fact that spin anisotropy remains robust against weak disorder in the AMMM phase. However, for strong disorder ($W\gtrsim 10t$), $G_{P}$ and $G_{AP}$ fully converge. This signals the complete suppression of spin anisotropy due to disorder scattering. Interestingly, both $G_{P}$ and $G_{AP}$ exhibit nonmonotonic behavior with increasing $W$, showing an enhancement over a finite range of $W$ before eventually merging and decreasing. This increase may be attributed to the fact that AM order reduces while the states are not fully localized. To quantify the sensitivity of tunneling to the AM bilayer configuration, we further compute the tunneling magnetoresistance $\text{TMR}=\frac{G_{P}-G_{AP}}{G_{AP}}$. As shown in Fig. 5(d), TMR depends on energy $E$. At $E=2t$, it decreases monotonically to zero, while at $E=0$, it first increases to a maximum before it eventually drops to zero. The latter case suggests that within a specific range of disorder strengths, the tunneling magnetoresistance can be anomalously enhanced.
@@
Conclusion and discussion.—In summary, we have shown that 2D metallic AMs persist as a marginal metal over a finite range of disorder strengths and undergo a KT metal-insulator transition as disorder increases further. We have also provided a phenomenological interpretation of this transition based on vortex-antivortex-like spin pairs. During the transition, the spin anisotropy of the AM gradually weakens, and the tunneling magnetoresistance becomes indistinguishable. These predictions are experimentally accessible and may help explain conflicting experiments on AM candidate materials.
@@
The key ingredient underlying our results is the interplay between AM order and magnetic disorder. Notably, neither this marginal metallic phase nor the KT-type metal-insulator transition appears in conventional antiferromagnets and nonmagnetic electron gases, where the electronic bands are spin-degenerate. Recently, many 2D AMs have been proposed and observed experimentally, including Mn2PSe with broken inversion symmetry Liu et al. (2025), $\mathrm{Rb_{1-\delta}V_{2}Te_{2}O}$ with layered structures Zhang et al. (2025); Jiang et al. (2025), ferroelectric systems Zhu et al. (2025b), monolayer Ma et al. (2021b) and twisted magnetic bilayers Liu et al. (2024b). Although our study focuses on representative $d$-wave AMs, we expect our conclusions to extend to other symmetry types, such as $g$- and $i$-wave AMs.
@@
We thank Lunhui Hu, K. Mæland, E. Petermann, and Jinsong Xu for helpful discussions. This work was supported by the DFG (SFB 1170 ToCoTronics), and the Würzburg-Dresden Cluster of Excellence ct.qmat, EXC 2147 (Project-Id 390858490). H.G. acknowledge support from the NSFC grant No. 12074022 and the BNL-CMP open research fund under Grant No. 2024BNL-CMPKF023. S.B.Z was supported by the start-up fund at HFNL, and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302801).
@@
In this section, we outline the transfer matrix method based on the lattice model to obtain the localization length of the system (MacKinnon and Kramer, 1983; Kramer and MacKinnon, 1993). Specifically, we consider a long ribbon geometry along $x$-direction with width $L_{y}=L$ and length $L_{x}\gg L$. In practice, we take $L_{x}=3\times 10^{6}$. Periodic boundary conditions are taken along $y$-direction. We partition the ribbon lattice to $L_{x}$ layers along $x$-direction. Focusing on three consecutive layers at $x=n-1$, $n$, and $n+1$, respectively, their wave function amplitudes can be connected by
@@
In Fig. 1 of the main text, we show the results for $t_{J}=0.3t$. Here, in Fig. S1(a), we further present the localization length for $t_{J}=2.0$ as a representative case of strong altermagnetic strength $t_{J}$. Clearly, the critical disorder strength $W_{c}$ shifts to higher values, reflecting the enhanced resilience of altermagnetic order. By fitting the scaling behavior to the form $\xi\propto\exp[b/\sqrt{W-W_{c}}]$ near $W_{c}$ [Fig. S1(b)], we find $W_{c}=11.27t\pm 0.33t$. We find that all data converge to a single curve under this scaling formula.
Figure S1: Kosterlitz-Thouless (KT) metal-insulator transition for $t_{J}=2.0t$. (a) Normalized localization length $\lambda/L$ as a function of $W$ for increasing $L$. (b) Single-parameter fitting of the correlation length $\xi$ near the KT-type transition, extracted from the data in (a). It gives the critical disorder strength $W_{c}=11.27t\pm 0.33t$. (c) Collapse of the data from (a) into a single curve under finite-size scaling. Other parameters are $t_{J}=2t$ and $\mu=-2t$.
@@
In this section, we describe in details how to obtain the disorder-averaged effective Hamiltonian $\langle H\rangle$. Consider the altermagnetic metal on a $L\times L$ square lattice with periodic boundary conditions in both $x$ and $y$ directions. For each disorder configuration, we find the retarded Green’s function $G^{R}({\bf r},{\bf r}^{\prime};\omega)$ in the tight-binding basis as
@@
It depends only on the position difference ${\bf r}-{\bf r}^{\prime}$. Next, we next Fourier transform the disorder-averaged Green’s function into momentum space,
@@
Finally, the disorder-average effective Hamiltonian is then given by
@@
The spin-resolved density of states then follows from
@@
In Fig. S2(a), we consider on-site nonmagnetic (spin-independent) disorder, i.e. $w_{i}^{0}\sigma_{0}$. Since this type of disorder does not mix opposite spins, the altermagnets (AMs) decouple into two independent spin sectors. From Fig. S2(a), we find that the normalized localization length always decreases with system size $L$. This indicates an immediate transition into insulating phases and the absence of Kosterlitz-Thouless (KT) type phase transition, consistent with conventional 2D electron behavior.
This behavior changes completely if we consider magnetic disorder, i.e., $(w_{i}^{x}\sigma_{x}+w_{i}^{y}\sigma_{y}+w_{i}^{z}\sigma_{z})$. Here, the diagonal type $w_{i}^{z}\sigma_{z}$ is included for comparison, although it does not flip spins. As shown in Fig. S2(b), $\Lambda=\lambda/L$ remains independent of $L$ within a wide range of disorder strengths, indicating the existence of a marginal metallic state. Beyond the critical value of disorder strength, $\Lambda$ decreases with $L$, indicating the KT-type metal-insulator phase transition. Comparing these results, we conclude that the magnetic scattering between opposite spins is essential for the occurrence of KT-type phase transitions in 2D AM metals.
@@
Appendix S4 Spectrum statistics in disordered altermagnetism from inverse participation ratio
In this section, we analyze the statistics from the eigenstates of the disordered AMs. To this end, we employ the inverse participation ratio (IPR) defined as (Li et al., 2017; Pixley et al., 2015)
$I_{n}=\sum_{i,\sigma}|\psi_{n}(i,\sigma)|^{4},$ (S4.1)
where the summation runs over all sites and spins. The disorder-averaged IPR $\langle I\rangle$ takes nearly zero values at weak disorder strength, consistent with the altermagnetic marginal metals discussed in the main text [Fig. S3(a)]. It becomes finite when the system enters an insulating phase.
The IPR scaling defines a fractal dimension $d_{2}$ via $\langle I\rangle\propto L^{-d_{2}}$. From Fig. S3(b), we find $d_{2}=1.84$ in the altermagnetic marginal metals, which is close to the value $d_{2}^{M}=2$ for ideal metals. For the insulating state at $W=14t$, we find $d_{2}=0.14$, consistent with the value $d_{2}^{I}=0$ for insulators. Between these two limits, a smooth transition is observed. Spectral rigidity, quantified by the spectrum compressibility
@@
also reflects the localization and delocalization behavior. Here, $\langle N\rangle$ is the disorder-averaged number of energy levels within an energy window. In 2D, $\chi$ is conjectured to be related to $d_{2}$ by $\chi=(2-d_{2})/4$. As shown in Fig. S3(c), we find $\chi=0.035$ at $W=3t$ and $\chi=0.44$ at $W=14t$, both in excellent agreement with the values inferred from $d_{2}$. Thus, the conjectured relation is confirmed in both the altermagnetic marginal metals and insulating states.

### v2

@@ Abstract
We introduce disorder into the two AM layers and calculate the tunneling conductances as functions of disorder strength $W$ [Fig. 5(c)]. We find that a substantial conductance difference between $G_{P}$ and $G_{AP}$ persists over a wide range of $W$, extending even to values $W>W_{c}$. This reflects the fact that spin anisotropy remains robust against weak disorder in the AMMM phase. However, for strong disorder ($W\gtrsim 10t$), $G_{P}$ and $G_{AP}$ fully converge. This signals the complete suppression of spin anisotropy due to disorder scattering. Interestingly, both $G_{P}$ and $G_{AP}$ exhibit nonmonotonic behavior with increasing $W$, showing an enhancement over a finite range of $W$ before eventually merging and decreasing. This increase may be attributed to the fact that AM order reduces while the states are not fully localized. To quantify the sensitivity of tunneling to the AM bilayer configuration, we further compute the tunneling magnetoconductance $\text{TMC}=\frac{G_{P}-G_{AP}}{G_{AP}}$. As shown in Fig. 5(d), TMC depends on energy $E$. At $E=2t$, it decreases monotonically to zero, while at $E=0$, it first increases to a maximum before it eventually drops to zero. The latter case suggests that within a specific range of disorder strengths, the tunneling magnetoconductance can be anomalously enhanced.
@@
Conclusion and discussion.—In summary, we have shown that 2D metallic AMs persist as a marginal metal over a finite range of disorder strengths and undergo a KT metal-insulator transition as disorder increases further. We have also provided a phenomenological interpretation of this transition based on spin vortex-antivortex pairs. During the transition, the spin anisotropy of the AM gradually weakens, and the tunneling magnetoconductance becomes indistinguishable. These predictions are experimentally accessible and may help to explain conflicting experiments on AM candidate materials.
@@
The key ingredient underlying our results is the interplay between AM order and magnetic disorder. Notably, neither this marginal metallic phase nor the KT-type metal-insulator transition appears in conventional antiferromagnets and nonmagnetic electron gases, where the electronic bands are spin-degenerate. Recently, many 2D AMs have been proposed and observed experimentally, including Mn2PSe with broken inversion symmetry Liu et al. (2025), $\mathrm{Rb_{1-\delta}V_{2}Te_{2}O}$ with layered structures Zhang et al. (2025a); Jiang et al. (2025), ferroelectric systems Zhu et al. (2025b), monolayer materials Ma et al. (2021b) and twisted magnetic bilayers Liu et al. (2024b). Although our study focuses on representative $d$-wave AMs, the conclusions hold for other symmetry types such as $g$-wave AMs and other lattice geometries such as the Lieb lattice. They remain valid even when the disorder is spatially correlated Li (2).
@@
We thank Fakher Assaad, Lunhui Hu, K. Mæland, E. Petermann, Jinsong Xu and Yan-Yang Zhang for helpful discussions. This work was supported by the DFG (SFB 1170 ToCoTronics), and the Würzburg-Dresden Cluster of Excellence ct.qmat, EXC 2147 (Project-Id 390858490). C.A.L. was also supported by the start-up fund at HFNL. H.G. acknowledges support from the NSFC Grant No. 12074022 and the BNL-CMP open research fund under Grant No. 2024BNL-CMPKF023. S.B.Z was supported by the start-up fund at HFNL, the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302801), and the National Natural Science Foundation of China (Grant No. 12488101).
Note added: Recently, we came across a related work that studies the influence of disorder on $d$-wave altermagnets with strong spin-orbit coupling and predicts a KT-type phase transition Zeng et al. (2025). However, our work substantially differs from Ref. Zeng et al. (2025) in that the KT-type transition in our case is solely induced by magnetic disorder in AMs in the absence of spin-orbit coupling and band topology.
@@
In this section, we outline the transfer matrix method based on the lattice model to obtain the localization length of the system (MacKinnon and Kramer, 1983; Kramer and MacKinnon, 1993). Specifically, we consider a long ribbon geometry along $x$-direction with width $L_{y}=L$ and length $L_{x}\gg L$. In practice, we take $L_{x}=3\times 10^{6}a$ in units of the lattice constant $a$. Periodic boundary conditions are taken along $y$-direction. We partition the ribbon lattice to $L_{x}$ layers along $x$-direction. Focusing on three consecutive layers at x=n1x=n-1, $n$, and $n+1$, respectively, their wave function amplitudes can be connected by
@@
In Fig. 1 of the main text, we show the results for $t_{J}=0.3t$. Here, in Fig. S1(a), we further present the localization length for $t_{J}=2t$ as a representative case of strong altermagnetic strength $t_{J}$. Clearly, the critical disorder strength $W_{c}$ shifts to higher values, reflecting the enhanced resilience of altermagnetic order. By fitting the scaling behavior to the form $\xi\propto\exp[b/\sqrt{W-W_{c}}]$ near $W_{c}$ [Fig. S1(b)], we find $W_{c}=11.27t\pm 0.33t$. All data converge to a single curve under this scaling formula.
Figure S1: Kosterlitz-Thouless (KT) metal-insulator transition for $t_{J}=2t$. (a) Normalized localization length $\lambda/L$ as a function of $W$ for increasing $L$. (b) Single-parameter fitting of the correlation length $\xi$ near the KT-type transition, extracted from the data in (a). It gives the critical disorder strength $W_{c}=11.27t\pm 0.33t$. (c) Collapse of the data from (a) into a single curve under finite-size scaling. Other parameters are $t_{J}=2t$ and $\mu=-2t$.
@@
In this section, we describe in detail how to obtain the disorder-averaged effective Hamiltonian $\langle H\rangle$. Consider the altermagnetic metal on a $L\times L$ square lattice with periodic boundary conditions in both $x$ and $y$ directions. For each disorder configuration, we find the retarded Green’s function $G^{R}({\bf r},{\bf r}^{\prime};\omega)$ in the tight-binding basis as
@@
It depends only on the position difference ${\bf r}-{\bf r}^{\prime}$. Next, we Fourier transform the disorder-averaged Green’s function into momentum space,
@@
Finally, the disorder-averaged effective Hamiltonian is then given by
@@
The spin-resolved density of states follows from
@@
In this section, we consider the scaling behavior of disordered altermagnets (AMs) under different disorder types. In Fig. S2(a), we consider on-site nonmagnetic (spin-independent) disorder, i.e., $w_{{\bf r}}^{0}\sigma_{0}$. Since this type of disorder does not mix opposite spins, the AMs decouple into two independent spin sectors. From Fig. S2(a), we find that the normalized localization length always decreases with system size $L$. This indicates an immediate transition into the insulating phase and the absence of the Kosterlitz-Thouless (KT) type phase transition, consistent with conventional 2D electron behavior. This behavior changes completely if we consider magnetic disorder of the type $(w_{{\bf r}}^{x}\sigma_{x}+w_{{\bf r}}^{y}\sigma_{y}+w_{{\bf r}}^{z}\sigma_{z})$. Here, the diagonal type $w_{{\bf r}}^{z}\sigma_{z}$ is included for comparison, although it does not flip spins. As shown in Fig. S2(b), $\Lambda=\lambda/L$ remains independent of $L$ within a wide range of disorder strengths, indicating the existence of a marginal metallic phase. Beyond the critical value of the disorder strength, $\Lambda$ decreases with $L$, indicating the KT-type metal-insulator transition. Comparing these results, we conclude that the magnetic scattering between opposite spins is essential for the occurrence of KT-type phase transitions in 2D AM metals.
@@
We further consider correlated disorder of Gaussian type. The disorder takes the form $V({\bf r})=V_{0}({\bf r})\sigma_{0}+V_{x}({\bf r})\sigma_{x}+V_{y}({\bf r})\sigma_{y}$, with zero mean $\langle V_{s}({\bf r})\rangle=0$ $(s=0,x,y)$ and spatial correlations $\langle V_{s}({\bf r})V_{s}({\bf r}^{\prime})\rangle=(W^{2}/12)\times\exp(-\left|{\bf r}-{\bf r}^{\prime}\right|^{2}/2\xi^{2})$, characterized by the correlation length $\xi$ and disorder strength $W$. A representative disorder configuration is shown in Fig. S3(a). To demonstrate the persistence of the marginal metallic phase and the associated KT-type phase transition, we calculate the scaling of the conductance as a function of $W$, see Fig. S3(b). The conductance curves merge together for different system sizes $L$ in the marginal metallic phase (for $W<W_{c}$) while they decrease with increasing $L$ in the insulating phase (for $W>W_{c}$) after a critical disorder strength $W_{c}$.
Figure S3: (a) A disorder configuration of the Gaussian-correlated disorder in real space with $\xi=1$ and $W=2t$. (b) Averaged conductance as a function of disorder strength $W$ for the correlated disorder. We take $2\times 10^{4}$ disorder configurations in the conductance scaling. Other parameters are $t_{J}=0.3t$ and $\mu=-2t$.
Appendix S4 Spectrum statistics from inverse participation ratio
In this section, we analyze the statistics of the eigenstates of the disordered AMs. To this end, we employ the inverse participation ratio (IPR) defined as (Li et al., 2017; Pixley et al., 2015)
$I_{n}=\sum_{{\bf r},\sigma}|\psi_{n}({\bf r},\sigma)|^{4},$ (S4.1)
where the summation runs over all sites and spins. The disorder-averaged IPR $\langle I\rangle$ takes nearly zero values at weak disorder strength, consistent with the altermagnetic marginal metals discussed in the main text [Fig. S4(a)]. It becomes finite when the system enters the insulating phase.
The IPR scaling defines a fractal dimension $d_{2}$ via $\langle I\rangle\propto L^{-d_{2}}$. From Fig. S4(b), we find $d_{2}=1.84$ in the marginal metal, which is close to the value $d_{2}^{M}=2$ for ideal metals. For the insulating state at $W=14t$, we find $d_{2}=0.14$, consistent with the value $d_{2}^{I}=0$ for insulators. Between these two limits, a smooth transition is observed. Spectral rigidity, quantified by the spectrum compressibility
@@
also reflects the localization and delocalization behavior. Here, $\langle N\rangle$ is the disorder-averaged number of energy levels within an energy window. In 2D, $\chi$ is conjectured to be related to $d_{2}$ by $\chi=(2-d_{2})/4$. As shown in Fig. S4(c), we find $\chi=0.035$ at $W=3t$ and $\chi=0.44$ at $W=14t$, both in excellent agreement with the values inferred from $d_{2}$. Thus, the conjectured relation is confirmed in both the marginal metallic and insulating phases.
@@
Appendix S5 $d$-wave altermagnets on the Lieb lattice
In this section, we consider $d$-wave AMs on a different lattice geometry, i.e., the Lieb lattice. There are three distinct sites in each unit cell, labeled as $A,B$, and $C$, respectively, as shown in Fig. S5(a). The $d$-wave AMs can be described by the Hamiltonian (Brekke et al., 2023)
$H_{\mathrm{Lieb}}=t\sum_{\langle ij\rangle\sigma}C_{i,\sigma}^{\dagger}C_{j,\sigma}-J\sum_{i,\sigma,\sigma^{\prime}}C_{i,\sigma}^{\dagger}{\bf S}_{i}\cdot{\bf\bm{\sigma}}_{\sigma\sigma^{\prime}}C_{i,\sigma^{\prime}}-\mu\sum_{i,\sigma}C_{i,\sigma}^{\dagger}C_{i,\sigma}+\varepsilon_{B}\sum_{i\in B,\sigma}C_{i,\sigma}^{\dagger}C_{i,\sigma},$ (S5.1)
where $t$ is the hopping amplitude, $J$ describes the exchange interaction strength between itinerant electrons and local spins ${\bf S}_{i}$, $\mu$ is the chemical potential, and $\varepsilon_{B}$ is the on-site energy at the nonmagnetic sites. Note that there is no spin-orbit coupling, thus the Hamiltonian is block diagonal in spin space $\sigma$. The Bloch Hamiltonian in momentum space for spin $\sigma$ reads
$H_{\sigma}({\bf k})=\left(\begin{array}[]{ccc}-\sigma JS-\mu&t(1+e^{ik_{x}})&0\\ t(1+e^{-ik_{x}})&-\mu+\varepsilon_{B}&t(1+e^{ik_{y}})\\ 0&t(1+e^{-ik_{y}})&+\sigma JS-\mu\end{array}\right).$ (S5.2)
We plot the energy spectrum along high-symmetry lines in Fig. S5(b). The energy bands exhibit anisotropic spin splitting, a characteristic feature of $d$-wave AMs.
Our results of disordered AMs can be generalized to the Lieb lattice. To illustrate this, we consider the same disorder of the form $(w_{i}^{0}\sigma_{0}+w_{i}^{x}\sigma_{x}+w_{i}^{y}\sigma_{y})$ at each lattice site and perform the scaling of the conductance of the $d$-wave AM on the Lieb lattice. We find that the conductance curves for different system sizes $L$ collapse when the disorder strength $W$ is below a critical value $W_{c}\simeq 3t$, signaling the emergence of a marginal metallic phase. For large $W$ above $W_{c}$, the conductance decreases with increasing $L$, consistent with an insulating phase. These results demonstrate the persistence of the marginal metallic phase and KT-type phase transition in the disordered $d$-wave AMs on the Lieb lattice.
Figure S5: (a) Altermagnet in a two-dimensional Lieb lattice, which contains three distinct sites in each unit cell, labeled by blue, gray, and red color respectively. (b) The electron bands of AMs on Lieb lattice along the high symmetry lines. (c) Averaged conductance as a function of disorder strength $W$. We average over $2\times 10^{3}$ disorder configurations. Other parameters for panels (b) and (c) are $JS=0.6t$, $\varepsilon_{B}=0$, and $\mu=-2t$.
Appendix S6 $g$-wave altermagnets and KT-type phase transition
In this section, we show that our results on $d$-wave AMs can be generalized to $g$-wave AMs in 2D. For illustration, we consider the minimum effective Hamiltonian for $g$-wave AMs (Ezawa, 2025)
$H_{g}({\bf k})=(-\mu+2t-t\cos k_{x}-t\cos k_{y})\sigma_{0}+J\sin k_{x}\sin k_{y}(\cos k_{x}-\cos k_{y})\sigma_{z},$ (S6.1)
where $t$ is the usual hopping amplitude, $J$ is the altermagnetic strength, $\mu$ is chemical potential, and $\sigma_{0}$ and $\sigma_{z}$ are the identity and Pauli matrices in spin space, respectively. The Fermi surface is shown in Fig. S6(a). It clearly exhibits the g-wave form of spin splitting. We consider the same onsite disorder of the form $(w_{{\bf r}}^{0}\sigma_{0}+w_{{\bf r}}^{y}\sigma_{y}+w_{{\bf r}}^{z}\sigma_{z})$ as in the main text. We perform the scaling calculation of the conductance for the $g$-wave AM, as shown in Fig. S6(b). The conductance curves converge for different system sizes $L$ when the disorder is smaller than a critical values $W_{c}\simeq 3.4t$, indicating the appearance of a marginal metallic phase. While the conductance decreases as the system size $L$ grows for disorder strength larger than $W_{c}$. These results demonstrate the presence of the marginal metallic phase and KT-type phase transition in the disordered $g$-wave AMs.
Figure S6: (a) Fermi surface of the $g$-wave AM at the Fermi energy, where the spin-up and spin-down states are indicated by different color and arrows. (b) Averaged conductance as a function of disorder strength $W$. We average over $2\times 10^{3}$ disorder configurations in the conductance scaling. Other parameters are $J=1.2t$ and $\mu=-0.6t$.
Appendix S7 Phenomenological picture based on spin vortices
In this section, we provide more details to support the phenomenological picture based on spin vortices in disordered AMs. We first briefly review the mechanism responsible for the KT phase transition. In the 2D XY model (and thin-film superfluids), the KT phase transition happens due to the formation of vortex-antivortex pairs (Kosterlitz and Thouless, 1973). At zero temperature, there are no vortices in the system. At low temperatures, bounded vortex-antivortex pairs are thermally activated. In this low-temperature regime, the system is in a quasi-ordered phase for the XY model and a superfluid phase for thin-film superfluids. As the temperature increases, the numbder of vortices increases and vortex-antivortex pairs become unbounded after a critical temperature, leading to the KT phase transition. At high temperatures above the transition, the system is in a disordered phase for the XY model and a normal fluid phase for thin-film superfluid.
The disorder-induced KT-type phase transition in 2D AMs can be explained along similar lines, with the disorder strength $W$ playing the role of temperature in the XY model or thin-film superfluid. In the 2D AMs with strongly anisotropic spin splitting, magnetic disorder of the form $(w_{{\bf r}}^{0}\sigma_{0}+w_{{\bf r}}^{x}\sigma_{x}+w_{{\bf r}}^{y}\sigma_{y})$ induces spin flips and generate pronounced in-plane magnetization components whose spatial texture forms a spin texture admitting vortices and antivortices (phase windings $\text{$\theta({\bf r})$=$\pm$}2\pi$). The typical in-plane spin textures for a vortex and anitvortex are schematically shown in Fig. S7(a). To substantiate this, we calculate numerically the real-space-resolved in-plane spin components at the Fermi level in presence of magnetic disorder:
$\bar{\sigma}_{s}({\bf r})=\sum_{n}\langle\Psi_{n}({\bf r})\left|\sigma_{s}\right|\Psi_{n}({\bf r})\rangle\rho(E_{n}),$ (S7.1)
where $\Psi_{n}({\bf r})$ is the eigenstate, $E_{n}$ the eigenvalue, and $\rho(E_{n})$ the density of states $\rho(E_{n})=\frac{\Gamma}{E_{n}^{2}+\Gamma^{2}}$ with $\Gamma$ a small band broadening.
We illustrate the results in Fig. S7(b-d). At $W=0$, there are no vortices in the system, just as in the 2D XY model and thin-film superfluid at $T=0$. Remarkably, in the weak-disorder regime, we observe bound vortex-antivortex pairs appear in the in-plane spin texture [see Fig. S7(b)]. As the disorder strength increases, the number of such pairs grows [see Figs. S7(c) and S7(d)], similar to thermally activated vortex pairs in the 2D XY model and thin-film superfluid. In this case, the system is in a marginal metallic phase, corresponding to the quasi-ordered phase in XY model or the superfluid phase in thin-film superfluids at low temperatures. In the strong-disorder regime, vortices significantly proliferate and unbind, that is to say, the phase field $\theta({\bf r})$ gets more and more distorted and in-plane spin correlations become short-ranged [see Fig. S7(d)]. In this case, the system is an insulator, corresponding to the disordered phase in the XY model or the normal fluid phase in thin-film superfluids at high temperatures. This progression, from bound pairs to vortex proliferation and vortex unbinding, provides a consistent, microscopic picture that underscores the KT transition induced by magnetic disorder in AMs. Note that no vortices appear for purely nonmagnetic disorder of $w_{{\bf r}}^{0}\sigma_{0}$ type or spin-polarized disorder of only $w_{{\bf r}}^{x}\sigma_{x}$ type and only $w_{{\bf r}}^{y}\sigma_{y}$ type. Accordingly, no KT-type phase transition occurs under these types of disorder. This further strengthens our physical picture based on the formation of in-plane spin vortices.
Figure S7: Formation of vortex and antivortex in the in-plane magnetization. (a) Sketch of vortex and antivortex in the in-plane spin components. (b) Spin texture at delocalized/marginal metal regime with disorder strength $W=0.5t$. Blue and yellow spots mark vortex and antivortex cores. (c) Same as (b) but at localized regime with $W=5t$, showing the proliferation of unbounded vortices and antivortices. (d) Averaged number of vortex pairs as a function of disorder strength $W$. Here, 50 disorder configurations are taken into the average. We use the parameters: $t_{J}=0.3t$, $\mu=-2t$, $\xi=2$, and $\Gamma=0.05t$ for illustration.